Use the given information to find the p-value. Also, use a 0.05 significance level and state the conclusion about the null hypothesis​ (reject the null hypothesis or fail to reject the null​ hypothesis).The test statistic in a​ right-tailed test is z=0.52.

A.

​0.6030; fail to reject the null hypothesis

B.

​0.3015; reject the null hypothesis

C.

​0.0195; reject the null hypothesis

D.

​0.3015; fail to reject the null hypothesis
Answer #1

p-valie = P(Z > 0.52) = 1 - P(Z < 0.52) = 1 - 0.6985 = 0.3015

since p-value is not less than significance level, we shoul not reject null hypothesis.

Option-D) 0.3015; fail to reject the null hypothesis